Two radar towers located 30 miles apart each detect an aircraft flying between them.

The angle of elevation measured by the first station is 32°. The angle measured by the second station is 18°.

Find the altitude of the aircraft.

   


1
Expert's answer
2021-06-15T08:38:30-0400

Given an arbitrary non-right triangle, we can drop an altitude, which we temporarily label "h,"

to create two right triangles.




From right triangle "ABD"


"x=\dfrac{h}{\tan \alpha}"

From right triangle "BCD"

"y=\dfrac{h}{\tan \beta}"


Then


"AC=x+y=\dfrac{h}{\tan \alpha}+\dfrac{h}{\tan \beta}=h(\dfrac{\tan \alpha+\tan \beta}{\tan \alpha\tan \beta})"


Solve for


"h=\dfrac{\tan \alpha\tan \beta}{\tan \alpha+\tan \beta}\cdot AC"

"h=\dfrac{\sin \alpha\sin \beta}{\sin \alpha\cos \beta+\sin \beta\cos \alpha}\cdot AC"

"h=\dfrac{\sin \alpha\sin \beta}{\sin (\alpha+\beta)}\cdot AC"

Given "\alpha=18\degree, \beta=32\degree, AC=32 \ miles."

Substitute


"h=\dfrac{\sin18\degree\sin 32\degree}{\sin (18\degree+32\degree)}\cdot 30\ mi\approx6.413 \ mi"

The altitude of the aircraft is 6.413 miles.